OpenStudy (anonymous):

It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? (Round answer to nearest tenth.) {Out______a0 hrs., back _______a1 hrs.}

OpenStudy (anonymous):

anything guys

hero (hero):

Okay, sorry. So to set this up use d = rt On the way there Distance = d1 Rate = 20mph time = t On the way back: Distance = d2 Rate = 30 mph Time = 12 - t d1 = 20t d2 = 30(12 - t) d1 = d2 20t = 30(12 - t) 20t = 360 - 30t (20 + 30)t = 360 50t = 360 t = 360/50 = 36/5 = (35 + 1)/5 = 7 + 1/5

hero (hero):

That's how many hours it took to get there. To get back do 12 - (7 + 1/5) 12 - 7 - 1/5 5 - 1/5 4 + 5/5 - 1/5 4 + 4/5 hrs to get back

hero (hero):

So 7 + 1/5 hours to get there 4 + 4/5 hours to get back

hero (hero):

To be more exact we can covert the fractions to minutes 1/5 of an hour is 12 minutes 4/5 of an hour is 48 minutes so... It took 7 hours 12 minutes to get there 4 hours and 48 minutes to get back